An employee who produces x units per hour earns an hourly wage of y = 0.80x + 11 (in dollars). Find the hourly wage for an employee who produces 19 units per hour.
Answer by checkley71(8403)

Y=.8X+11 OR Y=.8*19+11 OR Y=15.2+11 OR Y=$26.20 IS THE HOURLY SALARY FOR PRODUCING 19 UNITS PER HOUR.